OpenStudy (anonymous):

lovely little problem: given that n is an integer satisfying (n-3)^3 + n^3 = (n+3)^3 show that n must be even, and that n^2 must be a factor of 54 deduce that there is no such integer n

OpenStudy (anonymous):

given that m is an integer satisfying: (m-6)^3 + m^3 = (m+6)^3 deduce that m is even and hence show that there is no such integer m

OpenStudy (anonymous):

dain nevermind...

OpenStudy (anonymous):

has anyone solved it?

OpenStudy (anonymous):

does this have anything to do with the question or are these separate questions? (m-6)^3 + m^3 = (m+6)^3

OpenStudy (anonymous):

that is an extension to the question, if you have done the original question, try that one :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

subtract (n-3)^3 from both side results in: n^3 = 18n^2 + 54 the right side is clearly even so n^3 must be even... this implies n must be even.

OpenStudy (anonymous):

yeah :) i did that bit by contradiction, assuming n was odd

OpenStudy (anonymous):

how do you do that second part.. show n^2 is a factor of 54...

OpenStudy (anonymous):

you're almost there with the last post, each line is more of a hint, read as far as you like n^3 - 18n^2 = 54 n^2(n - 18) = 54 n is an integer, so n -18 is an integer hence n is a factor of 54

OpenStudy (anonymous):

oh... DUH!! :)

OpenStudy (anonymous):

i think i'm losing brain cells...

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (anonymous):

nice, although personally i prefer the method which doesnt require plotting a graph. has anyone attempted the second part?

OpenStudy (experimentx):

let x be an integer, 54 = 2 * 3 * 3 * 3 x^2 = +ve (x-18) ==> x > 18 ==> x^2 > 324 we have assumed x is an integer, (x-18) is also a positive integer x^2(x-18) > 324 <--- contradiction therefore .... x is not an integer!!

OpenStudy (kinggeorge):

Let's do the second question you posted. If we expand everything out, combine like terms, and set equal to 0, we get \[m^3=36m^2+432=2(18m^2+216)\]Since the right side is even, so is the left side. So , \(2|m^3\) so \(2|m \implies m\) is even. Like the other problem you posted, \(m^2(m-36)=432\) so it can be deduced that \(m^2|432\). Since \(432=2^4\cdot3^3\) the only possibilities for \(m^2\) are \(1, 4, 16, 9, 36, 144\). So \(m=\pm1, 2, 3, 4, 6, 12\). Now let's go to the equation \(m^2(m-36)=432\). Notice that \(m^2 >0\), and \(432>0\). Hence, \(m-36\) must be positive. However, none of our positive solutions allow that. Hence, there are no integral solutions.

OpenStudy (anonymous):

nice! i did it slightly differently, i'll post it

OpenStudy (anonymous):

(n-3)^3 + n^3 = (n+3)^3 assume n is odd we get: even + odd = even contradiction! hence n must be even expanding: n^3 - 18n^2 = 54 n^2(n - 18) = 54 n is an integer, so n -18 is an integer hence n^2 is a factor of 54

OpenStudy (anonymous):

since n is an even integer, let n = 2k n^2 = 4k^2 54 is not a multiple of 4 therefore we have a contradiction in other words there is no integer n: n^2(n-18) = 54

OpenStudy (anonymous):

for part 2 i used the same technique to show m was even, this time expanding gets: m^3-36m^2=432 m^2(m-36) = 432 notice 432 = 8 x 54 let m = 2r since its even 4r^2(2r - 36) = 432 8r^2(r-18) = 432 r^2(r-18) = 54 and we are done because we showed no integers satisfy this in part 1